# Efficient Depth Estimation for Unstable Stereo Camera Systems on AR Glasses

### v1

@@ 4.2 Homography Estimation Head
Unlike classical epipolar rectification, or the fast online rectification proposed by Argos [28], homography rectification only needs to transfer one image for alignment. However, one challenge is lost image information due to margins introduced in the rectified images.
Therefore, to eliminate the needs for homography rectification, we introduce a 2D rectification positional enconding (RPE) methodology that represents the position relationship between stereo images. This approach converts the homography matrix into positional encoding, which involves no information loss unlike homography rectification.
Following [26], we define the 2D positional encoding (PE) as:
@@
$i$$\in\{4k\leq i\leq 4k+3|i\in N,k\in N,k\in[0,C/4)\}$
where $q$ is a pixel located at coordinate $(x,y).~{}C$ is the number of channels of input activation. $i$ is the channel index. $f$ is the encoding frequency. The value of $f$ depends on the length of the sequence $len$ or the size of the input activation. Generally, it should satisfy $2\pi f>len$, and we default as $f=200$. Here $N$ means the set of natural numbers.
@@
Because both $PE(q_{l})$ and $RPE(q_{r})$ use the same coordinates, the same content in the stereo images gets a similar position encoding. Without RPE, stereo matching relies solely on images semantic similarity. With RPE, it also incorporates position similarity. In other words, if two image points are projections from the same world point, their similarity increases due to the additional positional information.
We present an example of the positional encoding results in Fig. 5 using colors representing regions with matched encoding values. The left image is obtained from Eq. 4 and the right is obtained from Eq. 5 under a given homography. We observe that the yellow pattern consistently appears on the right buckle of the leather strap, while the front edge of the power supply is always positioned between the red and pink patterns. The results show the efficacy of RPE in capturing alignment information in given stereo images or activations.
@@
Figure 6: Output comparison cross models and datasets
@@
Reconstructed from point clouds (need masks for invalid areas)
@@
AbsRel D1 RMSE AbsRel D1 RMSE AbsRel D1 RMSE AbsRel D1 RMSE AbsRel D1 RMSE
Dataset Sceneflow 0.172 0.71 8.2 0.129 0.42 5.9 0.287 0.89 68.8 0.102 0.23 5.3 0.091 0.35 4.3
ADT 0.199 0.36 611.9 0.135 0.52 563.3 0.176 0.55 546.1 0.133 0.34 320.1 0.094 0.25 273.6
DTU 0.147 0.38 315.9 0.148 0.40 224.8 0.339 0.88 628.5 0.122 0.61 536.9 0.101 0.42 216.1
Features CPU Latency (ms) 4906.2 50313.7 1184.2 811.0 625.4
GPU Latency (ms) Out of Memory Out of Memory 607.8 109.0 71.2
FLOPS(G) 26.32 190.14 95.80 7.27 7.28
#Params(M) 2.32 1.77 20.48 3.58 3.58
Table 2: The overall performance of the models. The datasets are used as inputs without any preprocessing. As indicated in Tab. 1, SceneFlow and ADT datasets consist of rectified stereo image inputs, while DTU contains unrectified stereo image inputs.
4.4 Structure of HomoDepth
In Fig. 4, we highlight training and inference flows using green and orange arrows, respectively. In the inference phase, the predicted homography matrix is passed to all the multi-head cost volumes, and the RPE is calculated inside it. The range of positional encoding is between 0 and 1, and it is added to the layer norm output. This ensures that the weights of input activations and positional encoding are balanced in the cost volume, allowing similarity to be assessed using both pattern and positional information.
During the training phase, the FC block and multi-head cost volume block are disconnected, resulting in two separate inputs and outputs in the model. The inputs include RGB stereo images and their homography matrix, while the outputs are the estimated homography and depth.
This structure is divided into three components: a common part (encoder), a depth estimation head, and a homography estimation head, as shown in Fig. 4. The shared encoder across two heads reduces overall computational costs and enables to perform two tasks with only a minor increase in latency for extra homography estimation head. However, training a multi-task model like this requires a specialized approach, which we discuss next.
4.5 Multi-task Model Training
To enable multi-task training of HomoDepth, we utilize the homoscedastic uncertainty [13] to train the model with the loss function adopted from Argos[28],
$L_{D}(y,\hat{y})$ $=smooth\_L_{1}(y,\hat{y})$ (6)
$+\sum_{l=1}^{5}smooth\_L_{1}(grad(y^{l}),grad(\hat{y}^{l}))$
where $smooth\_L_{1}$ is the SmoothL1Loss. $grad(y^{l})$ indicates the gradient of the depth map which is $2\times 2$ subsampled $l$ times. $y$ is the ground truth, and $\hat{y}$ is the prediction.
The loss function of homography estimation is formulated as follows:
$L_{H}(y,\hat{y})=||weight_{w}(y)-weight_{w}(\hat{y})||_{F}$ (7)
where
Utilizing the loss functions for homography and depth estimation. We formulate a unified total loss function as follows:
@@ 5 Evaluation
5.1 Datasets
We utilize Scene Flow Dataset (SceneFlow) [19], DTU Robot Image Data Sets (DTU) [10], and Aria Digital Twin (ADT) [20] datasets to evaluate our models and a state-of-the-art stereo depth estimation model for AR glasses, Argos [28]. We summarize the characteristics of the datasets in Tab. 1.
Aria Digital Twin (ADT) ADT consists of images captured by RGB and gray-scale fisheye cameras mounted on Aria AR glasses [5], a prototype AR glasses developed by Meta. Accordingly, the baseline and scenarios of the dataset are designed for AR applications. However, the ground truth depth in the dataset ignores human body in the images. To avoid disturbance from the property, we select 145 out of 236 scenarios that do not involve any human in the scene. Note that a scenario refers to a video that consists of many frames. We clarify details in Sec. 7.3.
ADT dataset provides one RGB image from the left and one gray-scale fisheye image from the right side for each scene due to the specification of the Aria glasses. Since both RGB and fisheye images are distorted, we apply calibration before feeding images to the models. Also, we replicate the calibrated grayscale images on channel dimension to match the shape of the RGB input tensor.
DTU Robot Image Data Sets (DTU) Dataset The DTU dataset is designed for multi-view stereo (MVS) reconstruction task. The dataset consists of images captured by a camera mounted on a robotic arm performing spherical scans, which move the robot arm from left to right with a slight downward tilt to the left back upon reaching the edge. We utilize images captured at neighboring positions as stereo image pairs for training. Compared to other MVS datasets, the DTU dataset is closer to the AR glasses scenarios for the following reasons: (1) Its scanning trajectory is close to the horizontal line, which can mimic the misalignment of stereo images. (2) The scanning intervals, or the baselines between adjacent images, are around 10 cm, which is similar to the cameras baseline on AR glasses.
We trained the models on each of these three datasets separately and compared the performance of the models. Specifically, since the DTU contains the extrinsic parameters of each frame, we leverage them in calculating the homography matrix over each stereo image inputs taken from two neighboring frames. We utilize the calculated homography matrix as the ground truth for the homography estimation in HomoDepth.
@@ 5.1 Datasets
Table 4: Ablation study comparing models with or without rectification preprocessing (PreP). DTU dataset represents a stereo system with a fixed focal length but variable relative position, while DTU_df represents a system with both variable focal length and position.
5.2 Implementation details
We implement MultiHeadDepth and HomoDepth in Pytorch [21] and train on Nvidia®  RTX 4090 32GB GPUs. We also run all the models involved in this section to obtain the test results. In experiment, we resize input images to $288\times 384\times 3$. The optimizer is Adam. In the early stages of model training, the base learning rate is 1e-4, and during the fine-tuning phase, it is adjusted to 4e-4.
5.3 Evaluation Metrics and Platform
We adopt AbsRel, D1, and RMSE as the evaluation metrics to measure model performance. All the metrics are ”Lower is Better” metrics, where smaller values represent higher model performance. We provide the detailed definition of each metric in Sec. 7.4.
Since our model’s eventual target is AR glasses, and we couldn’t find AR glasses with open API, we use three mobile platforms for latency evaluation as proxy. For the main model performance evaluation, we use a laptop with Intel® Core™ i7-12700H CPU and Nvidia® GeForce RTX 3070 Ti laptop GPU. We also measure the inference latency of our models on NVIDIA® Jetson Orin Nano™  Developer Kit [1] and a smartphone equipped with Snapdragon 8+ Gen 1 [2].
5.4 MultiHeadDepth Performance
We first evaluate the model performance on calibrated inputs. We compare the performance of MultiHeadDepth against four state-of-the-art stereo depth estimation models listed in Tab. 2. The results demonstrate that Argos and MultiHeadDepth are superior solutions for evaluated scenarios designed to model AR glasses. Comparing MultiHeadDepth against Argos, MultiHeadDepth provides 11.8-30.3% improvements in accuracy and 22.9-25.2% reduction in latency. The results show the effectiveness of our approach to optimize the cost volume blocks.
Impact of Quantization.
As quantization is commonly adopted optimization like Argos [28], we also evaluate quantized models and present the results in Tab. 3. In this study, we use AIMET [25] from Qualcomm to perform a post-training quantization. Although the overall accuracy decreases slightly by 3.1-14.7%, the latency is considerably reduced by 7.7-43.4% after quantization. MultiHeadDepth remains superior in all metrics after quantization.
5.5 HomoDepth Performance
In this session, we focus on the DTU scenarios that require preprocessing (rectification) and evaluate models with preprocessing and HomoDepth on such scenarios. We conduct the efficacy of RPE in the model through ablation studies of the end-to-end model (preprocessing + depth estimation). Utilizing the DTU dataset, we create a custom dataset, DTU_df, by scaling and cropping DTU images to simulate varying focal lengths, where ”df” stands for dynamic focal length. Thus, the DTU dataset represents a stereo system with a fixed focal length but variable relative position, while DTU_df represents a system with both variable focal length and relative position.
In Tab. 4 we evaluate four models with and without preprocessing. Although MultiHeadDepth achieves shorter inference latency, it still relies on preprocessing to achieve better accuracy, as compared in Tab. 2 and Tab. 4. In contrast, HomoDepth, despite having longer inference latency, does not require preprocessing. When the stereo cameras are misaligned, HomoDepth becomes a better solution. It not only delivers higher accuracy but also saves 30% of the total pipeline time.
5.6 Latency on Edge Devices
In this study, we report the latencies on two mobile/edge scale devices, NVIDIA® Jetson Orin Nano™  Developer Kit and Snapdragon 8+ Gen 1 Mobile Platform in Tab. 5. On the Orin Nano, the models are executed in eager mode. On the Snapdragon, the models are compiled using SNPE [22] and executed via ADB commands. The results may vary due to different deployment methods. We observe our models overall achieve better latency compared to Argos. Note that Argos results do not include preprocessing delay, while HomoDepth can accept unrectified data without preprocessing.
@@ 5.6 Latency on Edge Devices
Table 5: Latency of models in millisecond on NVIDIA® Jetson Orin Nano™  Developer Kit and Snapdragon 8+ Gen 1 Mobile Platform
@@ 6 Conclusion
Achieving both high model performance and low latency is a key toward depth estimation on AR glasses platforms. In this work, we identify preprocessing (rectification) and cost volume as the major optimization targets for low latency. By analyzing hardware-friendly and -unfriendly operations in existing cost volume, we develop an apporiximation of cosine similarity and employ a group-pointwise convolution that adopts the multi-head attention’s approach. For preprocessing, we introduce rectification positional encoding (RPE) and leverage homography matrix estimated by a head attached to the common encoder in our multi-head cost voulme blocks, which significantly reduces overall computational complexity.
@@
Our approaches are complementary to any stereo depth estimation systems with online rectification and/or cost volume. In addition to the broad applicability, our evaluation results indicate our approaches are effective in improving model performance as well as achieving low latency. Therefore, we believe our work is making a meaningful step forward in the field of stereo depth estimation systems.

### v2

@@ 4.2 Homography Estimation Head
Unlike classical epipolar rectification, or the fast online rectification proposed by Argos [31], homography rectification only needs to transfer one image for alignment. However, one challenge is the lost image information due to margins introduced in the rectified images.
Therefore, to eliminate the needs for homography rectification, we introduce a 2D rectification positional enconding (RPE) methodology that represents the position relationship between stereo images. This approach converts the homography matrix into positional encoding, which involves no information loss unlike homography rectification. Following [29], we define the 2D positional encoding (PE) as:
@@
$i$$\in\{4k\leq i\leq 4k+3|i\in\mathbb{N},k\in\mathbb{N},k\in[0,C/4)\}$
where $q$ is a pixel located at coordinate $(x,y).~{}C$ is the number of channels of input activation. $i$ is the channel index. $f$ is the encoding frequency. The value of $f$ depends on the length of the sequence $len$ or the size of the input activation. Generally, it should satisfy $2\pi f>len$, and we default as $f=200$. Here $\mathbb{N}$ means the set of natural numbers.
@@
Because both $PE(q_{l})$ and $RPE(q_{r})$ use the same coordinates, the same location in the real world leads to similar position encoding values in stereo images. Without RPE, stereo matching solely relies on images semantic similarity. However, RPE incorporates positional similarity, which enhances the similarity across two image points from the same world point utilizing the additional positional information.
We present an example of positional encoding in Fig. 6 using colors representing regions with matched encoding values. The left and right images are obtained under a given homography using Eq. 4 and Eq. 5, respectively. We observe that the yellow pattern consistently appears on the right buckle of the leather strap, while the front edge of the power supply is always positioned between the red and pink patterns. The results show the efficacy of RPE in capturing alignment information in given stereo images or activations.
@@
Rendered from point clouds (need masks for invalid areas)
@@
Figure 7: A comparison of output qualities across models and datasets
4.4 Structure of HomoDepth
In Fig. 5, we highlight training and inference flows using orange and green arrows, respectively. In the inference phase, the predicted homography matrix is passed to all the multi-head cost volumes, and the RPE is calculated inside them. The range of positional encoding is between 0 and 1, and it is added to the layer norm output. This ensures that the weights of input activations and positional encoding are balanced in the cost volume, allowing similarity to be assessed using both pattern and positional information. In the training phase, the FC block and multi-head cost volume block are disconnected, resulting in two separate inputs and outputs in the model. The inputs include RGB stereo images and their homography matrix, while the outputs are the estimated homography and depth.
Overall, the model structure is based on three components: a common part (encoder), a depth estimation head, and a homography estimation head, as shown in Fig. 5. The shared encoder across two heads reduces overall computational costs and enables to perform two tasks with only a minor increase in latency for extra homography estimation head. However, training a multi-task model like ours requires a specialized approach, which we discuss next.
4.5 Multi-task Model Training
To enable multi-task training of HomoDepth, we utilize the homoscedastic uncertainty [15] to train the model with two loss functions. The first loss function of depth estimation is adopted from Argos [31],
$L_{D}(y,\hat{y})$ $=SL_{1}(y,\hat{y})+\sum_{l=0}^{4}SL_{1}(\nabla y^{l},\nabla\hat{y}^{l})$ (6)
where $SL_{1}$ is the SmoothL1Loss [9]. $\nabla y^{l}$ indicates the gradient of the depth map which is $2\times 2$ subsampled $l$ times. $y$ is the ground truth, and $\hat{y}$ is the prediction.
The second loss function of homography estimation is,
$L_{H}(y,\hat{y})=||weight_{w}(y)-weight_{w}(\hat{y})||_{F}$ (7)
where
Utilizing the loss functions for homography and depth estimation, we formulate combined loss function as:
@@ 4.5 Multi-task Model Training
AbsRel D1 RMSE AbsRel D1 RMSE AbsRel D1 RMSE AbsRel D1 RMSE AbsRel D1 RMSE AbsRel D1 RMSE
Sceneflow 0.172 0.71 8.2 0.129 0.42 5.9 0.287 0.89 68.8 0.102 0.23 5.3 0.053 0.07 21.99 0.091 0.35 4.3
Middlebury 0.180 0.43 65.6 0.139 0.32 25.7 0.151 0.38 20.15 0.107 0.016 27.54 0.169 0.41 12.90 0.094 0.01 23.6
ADT 0.199 0.36 611.9 0.135 0.52 563.3 0.176 0.55 546.1 0.133 0.34 320.1 0.082 0.25 277.3 0.094 0.25 273.6
DTU 0.147 0.38 315.9 0.148 0.40 224.8 0.339 0.88 628.5 0.122 0.61 536.9 0.128 0.49 229.7 0.101 0.42 216.1
Table 2: The overall accuracy of the models. The datasets are used as inputs without any preprocessing. As indicated in Tab. 1, SceneFlow, Middleburry, and ADT datasets consist of rectified stereo image inputs, while DTU contains unrectified stereo image inputs. For Middleburry, following the practice in [31], models are trained with Sceneflow and tested on Middlebury 2014.
@@ 5 Evaluation
5.1 Datasets
We utilize SceneFlow [21], DTU Robot Image Datasets (DTU) [12], and Aria Digital Twin (ADT) [22] datasets to evaluate our models and a state-of-the-art stereo depth estimation model for AR glasses, Argos [31]. We summarize the characteristics of the datasets in Tab. 1.
Aria Digital Twin (ADT) ADT consists of images captured by RGB and gray-scale fisheye cameras mounted on Aria AR glasses [6], prototype AR glasses developed by Meta. Accordingly, the scenarios in the dataset are designed for AR applications. However, the ground truth depth ignores human body in the images. To avoid disturbance from the property, we select 145 out of 236 scenarios that do not involve any human in the scene. Due to Aria AR glasses’ hardware configuration, each scene in ADT consists of an RGB image from the left and a gray-scale fisheye image from the right. Since both RGB and fisheye images are distorted, we apply calibration before providing images to the models. Also, we replicate the calibrated grayscale images on channel dimension to match the shape of the RGB input tensor. We discuss more details about ADT in Appendix C.
DTU Robot Image Dataset (DTU) DTU dataset is designed for multi-view stereo (MVS) reconstruction task. The dataset consists of images captured by a camera mounted on a robotic arm performing spherical scans, which moves the robot arm from left to right with a slight downward tilt to the left back upon reaching the edge. We utilize images captured at adjacent positions as stereo image pairs for training. Compared to other MVS datasets, the DTU dataset is closer to the AR glasses scenarios since (1) its scanning trajectory is close to the horizontal line, which can mimic the misalignment of stereo images. (2) The scanning intervals between adjacent images (i.e., baselines), are $\approx$10 cm, which is similar as the baseline of AR glasses [6].
We train our model for the three datasets individually and evaluate the performance. Since the DTU contains the extrinsic parameters of each frame, we leverage them in calculating the homography matrix over each stereo image inputs taken from two adjacent frames. We utilize the calculated homography matrix as the ground truth for the homography estimation in HomoDepth. In the accuracy evaluation, we also evaluate our model on Middlebury 2014 [26].
Figure 8: The overall comparisons of models regarding # parameters, FLOPs, and latencies versus accuracy. For all coordinates, the closer to the bottom left, the better model performance. Here CPU indicates Core™ i7-12700H and GPU indicates Nvidia® GeForce RTX 3070 Ti laptop GPU.
@@ 5.1 Datasets
Table 4: Ablation study comparing models with or without rectification preprocessing (PreP). DTU dataset represents a stereo system with a fixed focal length but variable relative position, while DTU_df represents a system with both variable focal length and position. Sceneflow_persp is the simulation of glasses bending by perspective transformation.
@@
Table 5: Latency of models in millisecond on NVIDIA® Jetson Orin Nano™  Developer Kit and Snapdragon 8+ Gen 1 Platform.
5.2 Implementation details
We implement MultiHeadDepth and HomoDepth in Pytorch [23] and train on Nvidia RTX 4090 24GB GPU. We resize all input images in all experiments to $288\times 384\times 3$. We use Adam optimizer without schedulers. In the early stages of model training, the base learning rate (LR) is 1e-4, and we select the best epoch. Afterward, we fine-tune the model with an LR of 4e-4 and select the optimal weights. We set the batch size as 10, based on the GPU memory.
5.3 Evaluation Metrics and Platform
We adopt AbsRel, D1, and RMSE as the evaluation metrics to measure model performance. All the metrics are ”Lower is Better” metrics, where smaller values represent higher model performance. We provide the detailed definition of each metric in Appendix D.
Since our models’ eventual target is AR glasses, and there is no AR glasses with open API to the best of our knowledge, we use three mobile platforms for latency evaluation as proxy. For the main model performance evaluation, we use a laptop with Intel® Core™ i7-12700H CPU and Nvidia® GeForce RTX 3070 Ti laptop GPU. We also measure the inference latency of our models on Nvidia® Jetson Orin Nano™  Developer Kit [1] and a smartphone equipped with Snapdragon 8+ Gen 1 [2].
5.4 MultiHeadDepth Performance
We first evaluate the model performance on calibrated inputs. We compare the accuracy of MultiHeadDepth against SOTA stereo depth estimation models listed in Tab. 2. The overall results plotted in Fig. 8 demonstrate that Argos and MultiHeadDepth are superior solutions for evaluated scenarios designed to model AR glasses. Comparing MultiHeadDepth against Argos, MultiHeadDepth provides 11.8-30.3% improvements in accuracy and 22.9-25.2% reduction in latency. Compared to Selective-Stereo, MultiHeadDepth achieves competitive accuracy while using only 1.2% of the FLOPs required by Selective-Stereo. The results show the effectiveness of our approach to optimize the cost volume blocks.
Impact of Quantization As quantization is commonly adopted optimization like Argos [31], we also evaluate quantized models and present the results in Tab. 3. In this study, we use Aimet [28] from Qualcomm to perform a post-training quantization. Although the overall accuracy decreases slightly by 3.1-14.7%, the latency is considerably reduced by 7.7-43.4% after quantization. MultiHeadDepth remains superior in all metrics after quantization.
5.5 HomoDepth Performance
Since HomoDepth optimizes preprocessing, we focus on the DTU scenarios that require preprocessing (rectification) for evaluating HomoDepth. Utilizing the DTU dataset, we create a custom dataset, DTU_df, by scaling and cropping DTU images to simulate varying focal lengths (”df”: dynamic focal length). Thus, the DTU dataset represents a stereo system with a fixed focal length but variable relative position, while DTU_df represents a system with both variable focal length and relative position. Furthermore, we apply perspective transformation on Sceneflow (Sceneflow_persp) to simulate the misalignment of cameras.
In Tab. 5, we evaluate four models with and without preprocessing. Although MultiHeadDepth requires less inference latency than HomoDepth, it still relies on preprocessing to achieve better accuracy, as shown in Tab. 2 and Tab. 5. In contrast, HomoDepth does not require preprocessing. When the stereo cameras are misaligned, HomoDepth becomes a better solution, delivering higher accuracy with 30% less end-to-end latency.
5.6 Latency on Edge Devices
We report the latencies on two mobile/edge scale devices, Nvidia® Jetson Orin Nano™  Developer Kit and Snapdragon 8+ Gen 1 Mobile Platform in Tab. 5. On the Orin Nano, the models are executed in eager mode. On the Snapdragon, the models are compiled using SNPE [24] and executed via ADB commands. The results may vary due to different deployment methods. We observe our models overall achieve better latency compared to Argos. Note that Argos results do not include preprocessing delay, while HomoDepth can accept unrectified data without preprocessing.
@@ 6 Conclusion
Achieving both high model performance and low latency is a key toward depth estimation on AR glasses. In this work, we identify cost volume and preprocessing as the major optimization targets for low latency. For cost volume, we develop a hardware-friendly approximation of cosine similarity and employ a group-pointwise convolution. For preprocessing, we introduce rectification positional encoding (RPE) and leverage homography matrix estimated by a head attached to the common encoder in HomoDepth, which significantly reduces the computational complexity.
@@
Our approaches are complementary to any stereo depth estimation models with online rectification and/or cost volume. In addition to the broad applicability, our evaluation results show the effectiveness of our approaches in model performance as well as latency. Therefore, we believe our work is making a meaningful step forward in the field of stereo depth estimation systems for AR.
7 Acknowledgment
We thank Dilin Wang, Wei Ye, and Lianghzen Lai for their insightful feedback, Jialiang Wang for the guidance in modeling Argos, Meta for providing Aria AR glasses, and Rachid Karami for the assistance with the latency analysis.
Appendix A Computational Complexity of Multi-head Cost Volume
Suppose the dimension of an input activation is $CHW$, and the max disparity is $d$. The number of multiply-accumulate operations (#MAC) of the original cost volume is $3CHWd$ (please refer to Algorithm 1). The layer norm along channel dimension is defined as Eq. 10, whose #MAC is $3HWd$. Replacing the cosine similarity with the dot product, and adding the layer norm before the loop reduce the #MAC. The #MAC of multi-head cost volume is $2\times 3CHW+CHWd<3CHWd$. (The definition of parameters in Eq. 10 follow [4].)
$y=\frac{x-E[x]}{\sqrt{Var[x]+\epsilon}}\gamma+\beta$ (10)
Appendix B SIFT v.s. Conv Network
As [30] analyses, the computational complexity of SIFT for an $N\times N$ image with $n\times n$ tiles is
$\Theta\left(\frac{(n+x)^{2}}{p_{i}\,\Gamma_{0}}+\alpha\beta N^{2}x^{2}\Gamma_{1}+\frac{(\alpha\beta+\gamma)n^{2}\log x}{p_{o}\,\Gamma_{2}}\right)$ (11)
where x is the neighborhood of tiles. $N\gg n>x$ in most cases, we can simplify the complexity as $O(N^{2})$.
For convolutional networks, the computational complexity for an $N\times N$ input activation with $C$ channels in input and $C^{\prime}$ channels in output is,
$O(N^{2}CC^{\prime}k^{2})$ (12)
where $k$ is the convolution kernel size. $N\gg C>k$ in most cases, we can also simplify the complexity as $O(N^{2})$.
Based on the above analysis, we can conclude that: Supervised learning convolutional neural networks capable of the same task will not perform worse efficiency for all computer vision algorithms requiring key point matching. As the same, CNN is as efficient as, or even outperforms, classic algorithms in homography estimation tasks.
In practice, many optimizations for CNNs have been proposed, and CNN computations are more hardware-friendly. In contrast, [30] has been proven that without improvements in the input bandwidth, the power of multicore processing cannot be used efficiently for SIFT. Therefore, CNNs are generally a more efficient approach. Based on the report in [31] and our experiments, keypoint matching takes around 300ms on both smartphones and laptops. There is no significant speed up from smartphone to laptop, showing the limitations of keypoint matching.
Appendix C Dataset Setting
DTU setting:
Based on previous implementations and common practices [33], we selected the evaluation set as scans {1, 4, 9, 10, 11, 12, 13, 15, 23, 24, 29, 32, 33, 34, 48, 49, 62, 75, 77, 110, 114, 118}, validation set: scans {3, 5, 17, 21, 28, 35, 37, 38, 40, 43, 56, 59, 66, 67, 82, 86, 106, 117}, and the rest is training set.
Additionally, our network takes two images as input, but the depth map is aligned with the left image. This means that the left and right inputs cannot be interchanged. We need to match the stereo images to ensure that the relative position of the left input is indeed on the left side of the right input. The image match list is shown as Tab. 6.
ADT setting:
As mentioned, ATD ignores users’ bodies in the images when rendering ground truth depth maps which causes inconsistencies between the input images and the predicted results. We selected subsets of the scene where no other users were present. The selected subset can be obtained by this query link: https://explorer.projectaria.com/adt?q=%22is_multi_person+%3D%3D+false%22
L 1 2 3 4 5 7 8 9 10 11 12 12
R 2 3 4 5 6 6 7 8 9 10 11 13
L 13 14 15 16 17 18 19 21 22 23 24 25
R 14 15 16 17 18 19 20 20 21 22 23 24
L 26 27 28 29 29 30 31 32 33 34 35 36
R 25 26 27 28 30 31 32 33 34 35 36 37
L 38 40 41 42 43 44 45 45 46 47 48 49
R 39 39 40 41 42 43 44 44 45 46 47 48
Table 6: Input images pairs of DTU Dataset
Appendix D Metrics Definition
Here are the definitions of the metrics we used for evaluation:
$AbsRel$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\hat{y}_{i}-y_{i}}{y_{i}}$ (13)
$D1$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}(\frac{|\hat{y}_{i}-y_{i}|}{y_{i}}\leq 5\%)$ (14)
$RMSE$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}}$ (15)
where $N$ is the total number of pixels of the depth maps, $y_{i}$ is the value of pixels in ground truth map, and $\hat{y_{i}}$ is the value of pixels in prediction.
Appendix E Robustness Analysis
HomoDepth is a muti-task learning structure, and the depth estimation depends on the predicted homography matrix. Thus, the accuracy of homography estimation affects depth estimation. In this section, we address two questions: (1) How sensitive is the depth estimation to the homography estimation? (2) How stable is the homography estimation?
Homography Estimation Errors.
We analyze the error of HomoDepth during homography estimation. In practice, the error variance is as small as 0.003, as shown in Fig. 9(a). This indicates that the homography estimation of HomoDepth is highly accurate.
Sensitivity Study.
In HomoDepth, we inject noise $n\sim N(0,\sigma)$, where $\sigma\in[0,2]$, to the elements of estimated homography matrix before it is passed to the multi-head cost volume blocks. Then, we investigate the final depth estimation errors. The instability is quantified by examining the noise variance and corresponding changes in the smooth loss function Eq. 6 corresponding to depth estimation. As shown in Fig. 9(b), the standard deviation trends indicate that depth estimation remains stable when $\sigma<0.5$. The linear fitting demonstrates that the loss values increase as the noise variance grows.
(a) The statistics of the error in homography estimation. The curve represents the Gaussian fitting of the statistical results.
(b) The simulation of the influence of noise added to the homography matrix. The blue bars represent the smooth loss error ranges corresponding to noise variances $\sigma$, and the cross points represent outliers.
Figure 9: The robustness analyze of HomoDepth
Appendix F Application Scenarios